OpenStudy (waffelgirl12345):

will fan and medal! PLZ HELP!!! The gas tank in Sharon’s car can hold up to 16.5 gal of gas. About how many liters of gas can the tank hold? 1 L≈1.06 qt

OpenStudy (waffelgirl12345):

@sshayer

OpenStudy (waffelgirl12345):

@brainzonly

OpenStudy (mathmate):

1 gallon = 4 quarts 1 quart = 1.06 litre 16.5 gallons = \(16.5~gallons \times 4 \frac{quarts}{gallon}\) \(=16.5~gallons \times 4 \frac{quarts}{gallon}\times 1.06\frac{litres}{quart}\) \(=16.5\times 4\times 1.06 ~litres\) You'll need a calculator to complete your answer.

OpenStudy (waffelgirl12345):

so i use my calculator to do that last equation? 16.5*4*1.06?

OpenStudy (mathmate):

But please make sure you understand why and how the answer was obtained. From general knowledge, 1 gallon = 4 quarts, i.e. there are 4 quarts/gallon. Similarly, 1 quart=1.06 litres, or there are 1.06 litres/quart. What was written above will have all the units (gallons and quarts) cancelled except litre. So the final answer is 16.5*4*1.06 litres.

OpenStudy (sshayer):

1 litre=1.06 quart 1 quart=1.06 litre ? 1 galon=3.785 litres 16.5 galons=3.785*16.5=?